OpenStudy (anonymous):

solve for indicated letter h=4f, for f

OpenStudy (anonymous):

f = h/4

OpenStudy (anonymous):

so f is the answer

jimthompson5910 (jim_thompson5910):

f=h/4 is the answer

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

h/4=f